OpenStudy (anonymous):

What is the equation of the ellipse with foci (0, 4), (0, -4) and vertices (0, 8), (0, -8)? A. x squared divided by 16 minus y squared divided by 64 equals 1 B. x squared divided by 16 plus y squared divided by 64 equals 1 C. x squared divided by 48 plus y squared divided by 64 equals 1 D. x squared divided by 48 minus y squared divided by 64 equals 1 Ive narrowed it to B and C

OpenStudy (cj49):

(x-h)^2/b^2+(y-k)^2/a^2=1, a>b, (h,k) being the (x,y) coordinates of the center.

OpenStudy (cj49):

For given ellipse: center: (0,0) c=4 (from foci)

OpenStudy (anonymous):

I got B is that right? LOL

OpenStudy (cj49):

its C

OpenStudy (anonymous):

.-. ok... for sure?

OpenStudy (cj49):

it wud be x2/48+y2 /64

OpenStudy (anonymous):

Oh Oh ok i get it wha tbaout this one? What is the equation of the ellipse with co-vertices (-20, 0), (20, 0) and foci (0, -8), (0, 8)?

OpenStudy (anonymous):

(x^2/400) +(y^2+464)=1?

OpenStudy (cj49):

b2=400 c2=64 a2=b2+c2=400+64 x2/464 + y2/400=1